OpenStudy (anonymous):

How do you solve 7(4+2x) over or divided by (28+14x) =3

OpenStudy (anonymous):

\[\frac{7(4+2x)}{28+14x}=3\] \[7(4+2x)=3(28+14x)\] is a start

OpenStudy (anonymous):

then multiply out and add or subtract what you need to solve for \(x\)

OpenStudy (shamil98):

or. 7(4+2x)/7(4+2x) = 3 1 does not equal 3.

OpenStudy (shamil98):

no possible solutions *

OpenStudy (anonymous):

Thank you!